OpenStudy (anonymous):

what is the simplified form of 4x^2+16z+15/2z^2 -z -15

OpenStudy (anonymous):

15 + 4x^2 + 16z/-15 + 2z^2 + -z

OpenStudy (anonymous):

are you sure the variable x is correct, in 4x^2?

OpenStudy (anonymous):

its 4z^2 my bad

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

do you think you could solve it?

OpenStudy (anonymous):

15 + 16z + 4z^2/-15 + -1z + 2z^2 -1 + -16z + 2z^2 -1 - 16z + 2z^2 I think that is correct. Lemme check.

OpenStudy (anonymous):

the options are 2z+3/z-3 2z-3/z+3 2z+3/z+3 2z-3/z-3

OpenStudy (anonymous):

\[\frac{ 4z^2+16z+15 }{ 2z^2 -z-15 } = \frac{ (2z+5)(2z+3) }{ (2z+5)(z-3)} = \frac{2z+3}{ z-3} \]

OpenStudy (anonymous):

right yep,. here is the answer: 4x^2 + 16x + 15/2x^2 - x - 15 ( 2x + 3 )( 2x + 5 )/( x - 3 )( 2x + 5 ) 2x + 3/x - 3

OpenStudy (anonymous):

except instead of x use z